OpenStudy (anonymous):

solve for x and y in y=-5x-3

OpenStudy (ddcamp):

To solve for a variable, you have one of that variable alone (isolated) on one side of the equals sign. You can get a variable alone through addition, subtraction, multiplication, division etc.. Just remember that you have to do the same thing to both sides of the equation. What would you do to isolate x?

OpenStudy (anonymous):

5*x?

OpenStudy (ddcamp):

The equation has a 5*x already. What would you do to make that a 1*x (which is just x)?

OpenStudy (anonymous):

subtract 5? I don't know I stink at this kind of math

OpenStudy (ddcamp):

That's ok, it took us all a while to get it. Another way to think about it is "What's the opposite of multiplying by 5?"

OpenStudy (anonymous):

dividing by 5

OpenStudy (ddcamp):

Yup!

OpenStudy (ddcamp):

Basically, algebra is asking "How do I undo this" until you get to the place you want.

OpenStudy (ddcamp):

Another tip: Try undoing addition/subtraction before multiplication/division.

OpenStudy (ddcamp):

So, what would we want to do to isolate x in the original problem? y=-5x-3

OpenStudy (anonymous):

5/x

OpenStudy (ddcamp):

If we want to undo addition/subtraction first, we want to undo "subtract 3" first. What would this look like?

OpenStudy (anonymous):

y=-5*x?

OpenStudy (ddcamp):

OK, but you need to do the same thing to both sides: \[y = -5x -3 \\ y+3 = -5x -3 +3 \\ y+3 = -5x\]

OpenStudy (anonymous):

oh ok

OpenStudy (ddcamp):

Try doing the last step, by dividing both sides by -5

OpenStudy (anonymous):

so is y -5 or is x

OpenStudy (anonymous):

I'm confused

OpenStudy (ddcamp):

So we have negative five times x, right? To undo multiplication, we use division: